OpenStudy (anonymous):

8^2-x = 4^3x what is the value of x

OpenStudy (anonymous):

you have to balance the equation

OpenStudy (anonymous):

is this (8^2) - x = (4^3) x or is it 8^(2-x) = 4^(3x) ?

OpenStudy (anonymous):

sorry let me correct this 8^(2-x)=4^(3x)

OpenStudy (anonymous):

figured it out thanks anyways folks